OpenStudy (anonymous):

find the slope pf the line through each pair of points "show work" (3,-1) and (2,1)

OpenStudy (anonymous):

\[m = \frac{ y _{1}-y _{2} }{ x _{1}-x _{2} }\] \[m = \frac{ -1-1 }{ 3-2 }\] \[m = \frac{ -2 }{1}\] m = -2